OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half. y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (anonymous):

by putting them in a graphing calculator i know that the last two equations are the only ones that go through the point (6,-3) but im not sure how to find the slope

OpenStudy (anonymous):

one of them should be your answer though

OpenStudy (anonymous):

y=mx + b m is the slope and b is the y-int m=1/2, so y=1/2x+b to solve for b, use the given point given point: (6, -3) sub them into the equation -3= 1/2(6) + b solve for b, tell me what you got.